# Parameter-Efficient Augment Plugin for Class-Incremental Learning

### v1

@@ Abstract
Existing class-incremental learning (CIL) approaches based on replay or knowledge distillation are often constrained by forgetting or the stability-plasticity dilemma. Some expansion-based approaches could achieve higher accuracy. However, they always require significant parameter increases. In this paper, we propose a plugin extension paradigm termed the Deployment of extra LoRA Components (DLC) for non-pre-trained CIL scenarios. We treat the feature extractor trained through replay or distillation as a base model with rich knowledge. For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers. During inference, representations with task-specific residuals are aggregated to produce classification predictions. To mitigate interference from non-target LoRA plugins, we introduce a lightweight weighting unit. This unit learns to assign importance scores to different LoRA-tuned representations. Like downloadable contents in software, our method serves as a plug-and-play enhancement that efficiently extends the base methods. Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency. Moreover, it could surpass state-of-the-art methods under the fixed memory budget.
@@ 1 Introduction
In recent years, deep learning has achieved remarkable success across a variety of domains [jaimes2007multimodal, ye2019learning, chen2021large, chen2022learning, yang2024entaugment]. However, most models rely on a key assumption that the data distribution and learned classes remain fixed. In real-world scenarios, data distributions evolve continuously and new classes emerge over time [gomes2017survey], making static training paradigms impractical. Moreover, learning each new task typically relies on introducing additional task-specific sub-models, leading to inefficient model growth over time. Class-Incremental Learning (CIL) [zhou2024class, liangloss, yang2025supervised] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
Existing CIL methods  [zhou2024class] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches. Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [lopez2017gradient, rolnick2019experience, de2021rep]. However, the number of stored exemplars is typically far smaller than that of samples in the current task, creating a severe data imbalance that biases learning toward new classes and thus exacerbates catastrophic forgetting. Restriction-based methods [nguyen2017variational, ahn2019uncertainty] aim to mitigate catastrophic forgetting by constraining model updates during new task learning. A representative example is knowledge distillation [rebuffi2017icarl, simon2021learning], which preserves previous knowledge by aligning the current model’s output with those of its earlier versions. Such constraints introduce a trade-off: the model must balance adherence to past knowledge with adaptation to new data. This stability-plasticity dilemma fundamentally limits the achievable performance of restriction-based approaches.
@@
To address this, expansion-based CIL methods [mallya2018piggyback, serra2018overcoming, xu2018reinforced] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge and freezing previously learned parameters. While this isolates interference between tasks, the scalability limits as each new task requires its own feature extractor, causing parameters to grow linearly and incurring heavy storage costs. Recent pre-trained models-based methods [zhou2024continual] attempt to reduce this cost through lightweight fine-tuning modules. Yet, most PTM-based designs rely heavily on vision transformers with massive parameters (86M), and their fine-tuning modules can even surpass the size of conventional models. For instance, the introduction of a single Adapter module [zhou2024revisiting] incurs an parameter overhead (15M) greater than that of a complete ResNet-18 (11M). In addition, these PTM-based expansion methods face practical limitations, as PTMs are not universally available in practical applications. Therefore, expansion-based methods raise concerns despite strong performance. Their long-term sustainability and scalability in lifelong learning are questionable. This exposes a new quandary to CIL: How can we maintain high accuracy without a prohibitive linear increase in parameters as the number of tasks grows?
@@
To address this challenge, we propose the Deployment of extra LoRA Components (DLC). Rather than expanding the feature extractor model, DLC adopts a plug-and-play extension paradigm. Specifically, we attach task-specific Low-Rank Adaptation (LoRA) plugins to the deep layers of the base model. These plugins are updated exclusively during their corresponding task’s training phase, which occurs after the base model has been trained. This design ensures that the DLC plugins are trained independently, thus enabling their seamless and non-interfering integration into other methods. During inference, the base model loads all LoRA plugins to extract their feature representations, and they are aggregated into a composite one for final classification. However, non-target LoRA plugins that were not adequately updated on the dataset corresponding to the sample’s task may introduce noise or less informative residuals into the composite representation. So we propose a dedicated lightweight weighting unit before the classifier. This unit automatically assigns importance weights to different LoRA plugin outputs. DLC could enhance many non-expansion CIL approaches, such as knowledge distillation or replay methods. Experimental results on various benchmark datasets show that our method consistently improves the performance of replay- and distillation-based baselines. For instance, it could achieve up to 8% accuracy gains on ImageNet-100 using ResNet-18 and on CIFAR-100 using ResNet-32, while only increasing the total number of parameters by 4% and 11%, respectively. Furthermore, under the same memory budget, DLC outperforms existing state-of-the-art baselines. Our contributions are summarized as follows:
  • We propose a plugin extension framework DLC, which can be seamlessly integrated into diverse CIL methods without architectural modification by injecting task-specific residuals into deep layers via LoRA.
@@
  • We introduce a lightweight weighting unit for inference, which adaptively suppresses irrelevant LoRA activations according to their learned importance. This adaptive filtering mechanism effectively mitigates feature interference between old and new tasks, leading to more stable and accurate incremental classification.
@@
  • Extensive experiments show that DLC significantly boosts various CIL methods’ performance. Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA baselines.
@@ 2 Related Work
DLC can be viewed as a specialized form of expansion-based CIL that augments base methods with task-specific plugins. This section first introduces methods facing the stability-plasticity dilemma that are compatible with DLC, followed by a review of expansion approaches.
2.1 CIL with stability-plasticity dilemma
To mitigate catastrophic forgetting, many methods implement constrained parameter updates during model optimization [nguyen2017variational, bian2024make, cui2025few]. A prominent strategy in this category is knowledge distillation [simon2021learning, huang2024etag]. It protects the previously acquired knowledge through a modified distillation loss from the old to the new model. iCaRL [rebuffi2017icarl] mitigates forgetting by distilling the logits from the old model into the new model. Subsequent methods like BiC [wu2019large] and WA [zhao2020maintaining] further propose bias correction and weight alignment techniques to calibrate the classifier. Moreover, these methods preserve extra knowledge of prior classes by employing replay buffers [shaul2024teal, aghasanli2025prototype] to save exemplars in training. Furthermore, recent research has developed plug-and-play auxiliary loss functions [gao2025maintaining] that can be readily integrated into existing methods to enhance their accuracy. However, the fixed capacity of the feature extractor poses a fundamental constraint. The model is inherently limited in its ability to retain knowledge from past tasks and learn new tasks effectively, forcing it to seek a balance between stability and plasticity through a trade-off between classification and distillation/regularization losses, which ultimately caps its achievable accuracy.
2.2 Expansion-Based CIL
Such methods expand specialized parameters for model [mallya2018piggyback, serra2018overcoming, xu2018reinforced]. DER and TagFex [yan2021dynamically, zheng2025task] train an independent representation model on the new task for feature expansion, while the models of old tasks are frozen to retain the learned knowledge. MEMO [zhou2022model] splits the feature extractor into a generic shallow module and several task-specific deep modules. By increasing task-specific modules with fewer parameters, MEMO could save more memory space for storing replay samples than model-level expansion methods. Expansion-based methods achieve notably higher accuracy ceilings compared to fixed-capacity models. However, these methods result in a linear increase in the total number of parameters with tasks, imposing a significant storage overhead. Specifically, in the case of TagFex, the introduced attention module further increases the parameter overhead of the expansion. Even MEMO still expands over 70% of the extractor’s parameters for every new task. Emerging research leverages pre-trained models (PTMs) [zhou2024revisiting, mcdonnell2024ranpac, tan2024semantically] for CIL, mitigating catastrophic forgetting by leveraging knowledge encapsulated in PTMs. This has spurred the development of specialized parameter expansion methods designed for PTMs, which involve training lightweight fine-tuning modules like adapters [zhou2024expandable, sun2025mos]. However, such methods paradoxically result in higher overall parameter counts rather than a reduction. This is primarily due to the large parameter size of ViT backbones—significantly exceeding that of ResNet—combined with the non-negligible overhead of their tuning modules. An adapter introduces parameters equivalent to 15% of the base ViT’s total. Furthermore, PTM is not available in many CIL cases. These parameter-efficient module expansion methods depend on the rich prior knowledge embedded in PTMs and cannot be directly transferred to non-pre-trained models.
@@ 3.1 Problem Formulation of CIL
A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [de2021continual], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances. One class will appear only once on a given $D_{i}$ and will not be repeated. At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [rebuffi2017icarl] can be used for model training. For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively. The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
@@
Although the model’s parameters are dynamically updated throughout the incremental learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous tasks remains preserved. Replay ensures the model continues to minimize the empirical risk on a representative subset of past tasks, anchoring the parameters in regions consistent with previous data. Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model. Consequently, a model trained with the replay buffer and distillation retains significant knowledge about past classes, functioning as a capable but suboptimal ”generalist.” We posit that the knowledge stored in the model can be efficiently activated and tuned with minor task-specific parameters.
@@ 4 The Proposed Method
While the use of rehearsal and restriction enable the model to retain knowledge from previous tasks, an additional objective or imbalanced exemplars could create an optimization trade-off. This results in a suboptimal convergence point for the cross-entropy loss or past data, ultimately capping the model’s classification performance.
DLC proposes a distinctive paradigm, as shown in Fig. 2. We regard the model trained with a replay buffer and distillation as a foundation model, which stored knowledge across all tasks (albeit with suboptimal accuracy). Then refine it through residuals from each task-specific LoRA plugin to enhance accuracy. The details of the proposed DLC framework are elaborated in the following subsections.
4.1 LoRA for Convolutional Layers
A basic CIL model typically comprises a feature extractor $\phi(\boldsymbol{x})$ followed by a classifier $\boldsymbol{W}$. Most model expansion methods focus on expanding the $\phi(\boldsymbol{x})$ to achieve feature expansion. To maximize parameter efficiency, our approach avoids expanding the entire $\phi(\boldsymbol{x})$ or its deeper layers. Instead, we append parameter-efficient plugins after selected network layers within the $\phi(\boldsymbol{x})$ to adapt the representations produced by the base model. We specifically adopt Low-Rank Adaptation (LoRA) [hu2022lora] due to its minimal parameter size and its architectural flexibility.
The standard LoRA method was originally designed for linear layers. Given that most non-pre-trained CIL models are based on Convolutional Neural Networks (CNNs), it is necessary to change the original LoRA formulation to the convolutional operations fundamental to CNNs. For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation. By analogy to LoRA, ConvLoRA constructs a low-rank adaptation for a convolutional weight tensor via a sequence of two convolutional operations. The forward pass first processes the input with $\mathbf{A}$:

### v2

# Pushing the Limits of Distillation-Based Class-Incremental Learning via Lightweight Plugins

@@ Abstract
Existing replay and distillation-based class-incremental learning (CIL) methods are effective at retaining past knowledge but are still constrained by the stability–plasticity dilemma. Since their resulting models are learned over a sequence of incremental tasks, they encode rich representations and can be regarded as pre-trained bases. Building on this view, we propose a plug-in extension paradigm termed Deployment of LoRA Components (DLC) to enhance them. For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers. During inference, representations with task-specific residuals are aggregated to produce classification predictions. To mitigate interference from non-target LoRA plugins, we introduce a lightweight weighting unit. This unit learns to assign importance scores to different LoRA-tuned representations. Like downloadable content in software, DLC serves as a plug-and-play enhancement that efficiently extends the base methods. Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency. Under a fixed memory budget, methods equipped with DLC surpass state-of-the-art expansion-based methods.
@@ 1 Introduction
In recent years, deep learning has achieved remarkable success across a variety of domains [14, 36, 5, 4, 34]. However, most models rely on a key assumption that the data distribution and learned classes remain fixed. In real-world scenarios, data distributions evolve continuously and new classes emerge over time [11], making static training paradigms impractical. Moreover, learning each new task typically relies on additional task-specific sub-models, leading to inefficient model growth over time. Class-Incremental Learning (CIL) [42, 17, 35] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
Existing CIL methods  [42] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches. Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [18, 24, 8]. Other categories of methods likewise adopt rehearsal mechanisms to retain exemplars from past tasks, thereby further improving model accuracy. Expansion-based CIL methods [19, 25, 32] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge while freezing previously learned parameters. This strategy alleviates catastrophic forgetting and preserves plasticity on new tasks. However, rather than updating a single model incrementally, these methods expand the network and store, per task, an additional full [33] or partial (about 70%) [43] feature extractor, incurring substantial storage costs. For instance, with ResNet-18, each new task adds about 11.7MB of parameters, roughly equivalent to storing 3,998 images, which is about twice the replay buffer capacity under standard CIL settings. This overhead can limit their practicality for long task sequences, resource-constrained devices, and large-scale deployments.
@@
Restriction-based methods [21, 2] aim to mitigate catastrophic forgetting by constraining model updates on the new task. A representative example is knowledge distillation [23, 26], which preserves previous knowledge by aligning the current model’s output with its earlier versions. These methods perform CIL on a single model, leading to substantially lower storage overhead. As a result, these approaches are practically attractive in real-world settings. However, they inevitably introduce a trade-off [7]: the model must balance adherence to past knowledge with adaptation to new data. This stability-plasticity dilemma often becomes the key bottleneck for distillation-based approaches. Many enhancements [31, 38, 10] have been proposed to improve their accuracy, yet the core issue persists: distillation and new-task fitting are coupled objectives whose gradients can conflict, leaving performance governed by the same trade-off. A natural way to reduce this interference is to expand extra parameters for new tasks. However, scaling up the feature extractor is quickly parameter-prohibitive, and typical expansion-based schemes retrain task-specific extractors, breaking the shared single-model predictor that distillation relies on. This exposes a new quandary to CIL: How can we break the stability–plasticity dilemma by introducing a small set of task-specific auxiliary parameters, while remaining friendly to distillation-based methods?
@@
To address this challenge, we propose the Deployment of LoRA Components (DLC), a plug-and-play extension paradigm for CIL. We revisit parameter-efficient tuning modules in transfer learning beyond the frozen feature extractor assumption, and study them under the training dynamics of replay and distillation-based CIL. We attach task-specific lightweight plugins to deep layers of the base model and instantiate each plugin with LoRA in this work. These plugins provide additional knowledge by injecting residuals into the extracted features. For each incoming task, we first follow the underlying baseline to train the backbone while keeping all plugins frozen, and then update only the plugin associated with the current task. This decoupled update strategy trains each plugin independently, enabling DLC to be deployed into the baseline without interference. During inference, the model activates all plugins to extract their feature representations, which are aggregated for final classification. However, plugins that are not aligned with the input sample may contribute noisy or weakly informative residuals. We introduce a lightweight weighting unit before the classifier, which automatically assigns importance weights to different plugin outputs. Experimental results on multiple benchmarks show that DLC consistently improves replay and distillation-based baselines. For example, DLC achieves up to 8% accuracy gains on ImageNet-100 with ResNet-18 and on CIFAR-100 with ResNet-32, while increasing the total parameter count by only 4% to 30%. Under the same memory budget, DLC also outperforms existing state-of-the-art CIL baselines. Our contributions are summarized as follows:
  • We propose DLC, a plug-and-play plugin extension framework that enables lightweight task-specific plugin residual injection in continually updated backbones, anchored by the consistent knowledge preserved through replay and distillation.
@@
  • We introduce a lightweight weighting unit that adaptively suppresses irrelevant plugin activations based on learned importance, reducing feature interference across tasks and leading to more stable and accurate incremental classification.
@@
  • Extensive experiments show that DLC significantly boosts various CIL methods’ performance. Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA CIL methods.
@@ 2 Related Work
2.1 Class Incremental Learning
Class-incremental learning primarily aims to prevent the model from forgetting old classes. Existing methods tackle this challenge from multiple perspectives, including rehearsal [16, 1] that replays representative exemplars from previous classes, regularization [3] or distillation [12] that limit overly aggressive updates on new data, expansion-based designs that train and retain task-specific full [33] or partial [43] feature extractors, and further integrate contrastive objectives [39] to stabilize representations. In addition, some methods leverage transfer learning with Transformers pretrained on large-scale datasets to provide stronger representations [40, 27]. These methods often exhibit an accuracy–memory trade-off: approaches with higher recognition accuracy typically require storing more replay samples or allocating and maintaining additional model parameters through expansion or larger backbones.
2.2 Distillation-Based CIL
Distillation-based CIL trains the model on new tasks while regularizing it to match the predictions of the previous version. A wide range of distillation-based approaches have been explored, including combining distillation with replay [23], incorporating bias correction [31] or weight alignment [38], momentum-based distillation [20], and plug-and-play auxiliary objectives [10]. Distillation remains appealing as it introduces minimal overhead, can leverage unlabeled incoming samples, and fits a simple single-model training loop, making it broadly applicable in practical settings such as online [30], imbalanced [22], and few-shot [37] CIL. However, the feature extractor’s fixed capacity imposes a fundamental constraint: the model must balance stability and plasticity by trading off classification and distillation losses, ultimately capping achievable accuracy.
@@ 3.1 Problem Formulation of CIL
A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [7], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances. One class will appear only once on a given $D_{i}$ and will not be repeated. At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [23] can be used for model training. For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively. The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
@@
Although the model’s parameters are dynamically updated throughout the incremental learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous tasks remains preserved. Replay ensures the model continues to minimize the empirical risk on a representative subset of past tasks, anchoring the parameters in regions consistent with previous data. Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model. Consequently, a model trained with replay and distillation preserves substantial knowledge of past classes, resulting in a single unified model that is broadly effective but suboptimal. We posit that the knowledge stored in the model can be efficiently activated with minor task-specific parameters.
@@ 4 The Proposed Method
DLC proposes a plugin-level residual injection paradigm, as illustrated in Fig. 2. We treat a replay and distillation-trained network as a base model that accumulates knowledge across tasks but remains suboptimal in accuracy, and then augment it with task-specific plugins that inject residual knowledge into the extracted features to improve performance. In this work, each plugin is instantiated with a LoRA-style low-rank parameterization, but is used as a lightweight task extension rather than a conventional fine-tuning update. The proposed DLC is detailed in the following subsections.
4.1 Task-Specific Plugins Deployment
Parameter-efficient tuning is typically used in transfer learning by adapting a feature extractor with frozen parameters, where layer inputs remain largely stationary. Here, we revisit it beyond this static-backbone assumption and study it under replay- and distillation-based class-incremental learning, where old-task outputs are constrained, and intermediate feature drift is therefore controlled during continual updates. Concretely, after training on task $t$, the output deviation at layer $\ell$ is bounded by $K_{\ell}\Gamma_{t}$ (see Appendix), where $K_{\ell}$ depends on hyperparameters such as network structure and distillation temperature, and $\Gamma_{t}$ reflects the distribution discrepancy induced by the chosen replay and distillation strategy. This bound directly limits the worst-case feature drift on the old-task distribution, enabling stable optimization of lightweight auxiliary parameters in this dynamic regime.
We introduce a plugin extension design and maintain an independent plugin set for each task. Specifically, for each task $t$, we create a dedicated set of $k$ plugins for $\phi(\boldsymbol{x})$, denoted as $L_{t}$. The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$. In DLC, plugins are deployed in the final layers of $\phi(\boldsymbol{x})$. Taking a CNN-based architecture as an example, the first plugin in $L_{t}$ is attached to the last convolutional layer of $\phi(\boldsymbol{x})$, and the remaining plugins are placed on progressively earlier layers moving backward through the network. This design is motivated by two considerations. First, deeper features are more semantically meaningful and are directly used for classification, so injecting task-specific residuals at this level can more effectively enhance inter-class discriminability. Second, it is more compatible with distillation-based objectives. If plugins are deployed in shallow layers, residual changes in low-level features would propagate through the network and amplify logit deviations, increasing the optimization difficulty of distillation losses that rely on output alignment between old and new models.
In DLC, the plugins and $\phi(\boldsymbol{x})$ undergo a decoupled two-phase training procedure. Specifically, when a new task $t$ arrives, we first train the base model following the underlying baseline procedure, updating only $\phi(\boldsymbol{x})$ and the classifier $\boldsymbol{W}$ while keeping all plugins frozen. After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and $L_{t}$ is trained with the same data. Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge. The training objective for all plugins incorporates the cross-entropy loss and an auxiliary loss [33] commonly used in expansion-style methods, formulated as:
$\mathcal{L}_{\text{plg}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (5)
The two-phase procedure ensures that plugins are trained without interference from distillation losses, enabling specialized learning for the target task. Since $\phi(\boldsymbol{x})$ remains frozen during plugin updates, plugin training is stable, and it also prevents the baseline training of $\phi(\boldsymbol{x})$ from being affected by plugin training, preserving the plug-and-play property of DLC.
During inference, all plugins are activated to extract their respective feature representations. These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble. For a linear classifier $\boldsymbol{W}$, the classification process can be formulated as:
where $T$ denotes the maximum index of known tasks. Notably, we do not need to run the full extractor $T$ times: each plugin injects a residual only at its deployment layer, so the backbone activations up to the earliest deployed layer are shared and computed once, avoiding redundant computation and improving inference efficiency.
4.2 LoRA-Style Residual Plugin
To instantiate each task-specific plugin with minimal parameters, we adopt a LoRA [13]-style low-rank residual parameterization. This design is well-suited for plugin deployment: the rank $r$ provides a simple capacity knob under a fixed budget, the same low-rank residual form applies to both linear and convolutional operators, and the residual structure supports our decoupled training and inference aggregation since plugins can be independently trained, frozen, and combined without modifying the backbone. We emphasize that the backbone weights are kept intact throughout, and the low-rank component is introduced as an auxiliary module that adds a task-specific residual to the feature. That is, the LoRA plugin does not replace or factorize the original weights, but injects additional task knowledge via an additive low-rank branch, which aligns with our plug-and-play deployment setting.
Since most CIL pipelines train CNN-based backbones from scratch, we instantiate the plugin for convolutional operators using a LoRA-style low-rank residual branch (implemented as two convolutions). For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation. The LoRA plugin first processes the input with $\mathbf{A}$: